OpenStudy (sasogeek):

differentiate the function \[\ y=frac{1}{(1-5x)^2} .... I'm getting \[\10(1-5x)^-3 \] my "book" says \[\ -10(1-5x)^-3 \] if the book is wrong again... i'll stop using it! :@ but if i'm wrong, tell me what i did wrong cos there aren't steps in the book, it's just a question and a solution

OpenStudy (sasogeek):

looks like my latex is messed up there... sorry but i'm not good at it

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{d}{dx}(1-5x)^{-2}=-2(1-5x)^{-3}*\frac{d}{dx}(1-5x)=-2(-5)(1-5x)^{-3}\] \[\implies \frac{dy}{dx}=10(1-5x)^{-3}\]

OpenStudy (sasogeek):

that's it! i'm done with this book, thanks malevolence, i'm going to look for a different source for questions :)

myininaya (myininaya):

You always write pretty things male

OpenStudy (anonymous):

Thanks :) I like my math haha

OpenStudy (phi):

Use wolfram to check your answers.